### v3

# ThinkNote: Enhancing Knowledge Integration and Utilization of Large Language Models via Constructivist Cognition Modeling

Abstract
Large Language Models (LLMs) have demonstrated strong performance across a wide range of NLP tasks. However, they often exhibit suboptimal behaviors and inconsistencies when exposed to unfamiliar external information, underscoring their limitations in effectively leveraging such knowledge. Inspired by constructivist learning theory, we propose ThinkNote, a novel framework that enhances the external knowledge utilization of LLMs through a two-stage constructivist cognitive modeling process. Specifically, ThinkNote performs knowledge assimilation to align new information with the model’s parametric memory, forming a coherent internal representation. It then applies thought accommodation to adapt internal reasoning, thereby promoting more consistent and reliable outputs. Extensive experimental results demonstrate that ThinkNote achieves a 10% improvement over strong baseline methods on various question-answering benchmarks. Further analysis indicates that ThinkNote effectively integrates and utilizes external knowledge to help LLMs generate accurate responses and improves their self-consistency. All data and codes are available at https://github.com/OpenMatch/ThinkNote.
ThinkNote: Enhancing Knowledge Integration and Utilization of Large Language Models via Constructivist Cognition Modeling
Zhipeng Xu1, Zhenghao Liu1thanks: indicates corresponding author., Yukun Yan2, Shuo Wang2, Shi Yu2 Zheni Zeng2, Chaojun Xiao2, Zhiyuan Liu2, Ge Yu1 and Chenyan Xiong3 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
1 Introduction
Large Language Models (LLMs), e.g. GPT-4 OpenAI (2023) and LLaMA Touvron et al. (2023), have shown strong emergent abilities and achieved convincing performance in various NLP tasks Wei et al. (2022a); Zhao et al. (2023a). Despite these advancements, LLMs often face challenges in effectively leveraging external knowledge to generate accurate responses, especially when encountering information that is noisy or incomplete Liu et al. (2023); Xie et al. (2024); Asai et al. (2023). Such limitations lead to response inconsistencies and hallucinations, constraining their application in knowledge-intensive scenarios Ji et al. (2023); Xu et al. (2024a); Wang et al. (2023).
Recent research investigates strategies for enhancing the integration and utilization of external knowledge in LLMs. These efforts aim to develop specialized modules that refine, reflect on, or summarize external knowledge before integration Yu et al. (2023a); Xu et al. (2023) or filter the external knowledge based on its relevance to the current context Peng et al. (2023); Zhao et al. (2023b); Asai et al. (2023). While these methods show their effectiveness in helping LLMs acquire new knowledge Ji et al. (2023); Wei et al. (2024), they still treat LLMs as passive recipients of information Steffe and Gale (1995); Larochelle et al. (1998) and fail to address the underlying cognitive limitations that lead to insufficient knowledge utilization Yu et al. (2025). This limitation leads to incorrect responses and even causes degenerated responses on certain tasks Foulds et al. (2024); Shuster et al. (2021).
To address these limitations, we propose a shift from treating LLMs as passive consumers of knowledge to modeling them as active knowledge constructors, drawing inspiration from constructivist cognitive theories Steffe and Gale (1995); Papert (1980); Piaget and Inhelder (1972). Constructivism emphasizes that knowledge is not simply absorbed from the environment, but is actively built through experience, reflection, and assimilation of new information into existing mental frameworks von Glasersfeld (1984). Under this paradigm, external knowledge is not directly fused into the input context of LLMs but transformed and recontextualized based on the model’s current understanding of the task or question Schunk (2012). Such a cognitively grounded design encourages models to resolve inconsistencies, discard irrelevant or misleading information, and synthesize new insights, similar to how humans build reliable understanding from noisy data Bruner (1990).
In this paper, we propose ThinkNote, a cognition-inspired framework to activate LLMs to effectively integrate and utilize external knowledge by mimicking the cognitive processes of assimilation and accommodation from constructivism Steffe and Gale (1995). Specifically, ThinkNote incorporates knowledge assimilation to transform external information into coherent insights and understanding, extending beyond query-specific notes Yu et al. (2023a) to enrich the parametric knowledge of LLMs. The thought accommodation then updates the parametric memory of LLMs by incorporating the assimilated knowledge to refine and enhance the internal reasoning chain of LLMs. This two-step cognitive process is prompted to mimic different learning behaviors of humans, where we configure ThinkNote using four instructions to process external knowledge by associating familiar knowledge, anchoring unfamiliar information, extracting logical reasoning, or identifying counterfactuals.
Our experiments in various knowledge-intensive question-answering tasks demonstrate the effectiveness of ThinkNote. Compared to utilizing the external knowledge directly, ThinkNote achieves a 10% improvement regardless of the scale of the LLMs used as backbone models. Our further analysis shows that ThinkNote causes information accumulation Fan et al. (2025) on specific tokens, suggesting that ThinkNote can effectively utilize critical elements of external information in generating responses.
2 Related Work
Large Language Models (LLMs) consistently face stability challenges when utilizing external knowledge Zhao et al. (2023b); Asai et al. (2023). While LLMs are exposed to extensive corpora during training and demonstrate strong generalization across various tasks Zhao et al. (2023a); Touvron et al. (2023), their performance declines when encountering unfamiliar, ambiguous, or contradictory information Ji et al. (2023); Liu et al. (2023); Foulds et al. (2024). These limitations illustrate a notable gap between the availability of external knowledge and the model’s capacity to effectively integrate and apply it, especially evident in knowledge-intensive tasks Yu et al. (2023b); Shi et al. (2023); Jiang et al. (2023).
To mitigate these limitations, some methods attempt to enhance knowledge integration by optimizing the interface between external information and the model’s input context Cuconasu et al. (2024); Gao et al. (2023b). These methods typically filter or reformat noisy inputs through knowledge summarization Xu et al. (2023), reflection Asai et al. (2023), or multi-step verification Yu et al. (2023a); Peng et al. (2023). While these approaches have advanced the ability of LLMs to integrate and utilize external information, they often push the burden of knowledge alignment to external modules Wang et al. (2024), where the LLMs remain conceptualized as passive processors, failing to reorganize or evaluate new information based on their current understanding Madaan et al. (2023).
More recent efforts explore internal modifications to how LLMs represent, revise, or adapt knowledge. These include work on memory editing Meng et al. (2022); Trivedi et al. (2022), belief tracking Wilie et al. (2024), and metacognitive strategies such as self-verification and response calibration Madaan et al. (2023); Liu et al. (2023). While promising, these approaches are often fragmented and tool-like: they provide mechanisms to intervene in model behavior without modeling the interaction between the model and external knowledge Collins et al. (2024). There is still limited consensus on what constitutes effective knowledge integration at the cognitive level, and most current methods operate under engineering-driven constraints rather than principles drawn from learning theory or epistemology Banino et al. (2023); Floridi and Nobre (2024); Van Rooij et al. (2024).
3 Methodology
In this section, we introduce our ThinkNote framework, grounded in constructivist cognitive theory Steffe and Gale (1995), which aims to emulate the cognitive process of knowledge learning employed by human learners to interact with external knowledge. We first describe the overall workflow of ThinkNote framework (Sec. 3.1) and then elaborate on the specific configuration strategies used to regulate learning behavior (Sec. 3.2).
Figure 1: The Overview of the ThinkNote Workflow.
3.1 Enhancing Knowledge Integration and Utilization in LLMs via ThinkNote
As shown in Figure 1, ThinkNote is a cognition-inspired framework designed to enhance the ability of LLMs to integrate and utilize external knowledge to generate accurate responses, which draws on the cognitive principles of constructivist learning theory Steffe and Gale (1995). This framework simulates a two-stage cognitive process: ThinkNote first conducts knowledge assimilation to align new information with the model’s parametric memory, and then applies thought accommodation to adapt internal reasoning, thereby promoting more consistent and reliable outputs.
Knowledge Assimilation. Humans activate relevant memory (which is referred to as schema in constructivist cognitive theory) to interpret the meaning and function of external information and construct new understanding, which is called knowledge assimilation Steffe and Gale (1995); Moreno and Mayer (1999). The knowledge assimilation stage in ThinkNote simulates this process, driving the LLMs to align and integrate external knowledge with their internal parametric memory and point out the function of knowledge for problem-solving. Specifically, given a query $q$ and a set of external knowledge $\mathcal{D}$, ThinkNote guides the foundational model $\mathcal{M}$ to generates an assimilated rationale $\mathcal{T}$, which can be expressed as:
$\mathcal{T}=\mathcal{M}\left(\mathcal{I}_{\text{ka}},q,\mathcal{D}\right),$ (1)
where $\mathcal{I}_{ka}$ indicates the instruction to mimic different learning behaviors. The assimilated knowledge $\mathcal{T}$ extends beyond knowledge summarization Yu et al. (2023a) and serves to contextualize the knowledge within the problem scenarios, enabling the model to not only recall relevant facts but also explain how and why these facts matter.
Thought Accommodation. The thought accommodation refers to the adjustment of internal memory when new information challenges prior understanding in constructivist theory Steffe and Gale (1995); Larochelle et al. (1998). In ThinkNote, this process is simulated by guiding the model $\mathcal{M}$ to revise its initial reasoning based on assimilated knowledge $\mathcal{T}$. Given a query $q$, ThinkNote first conducts self-inquiry to generate a preliminary chain-of-thought $\mathcal{R}$ using its parametric memory:
$\mathcal{R}=\mathcal{M}\left(\mathcal{I}_{\text{CoT}},q\right).$ (2)
Here $\mathcal{I}_{\text{CoT}}$ triggers a chain-of-thought reasoning grounded in prior knowledge, which may lack context or accuracy due to outdated or incomplete parametric memory. To address this, ThinkNote guides the model $\mathcal{M}$ to revise the chain-of-thought $\mathcal{R}$ using the assimilated knowledge $\mathcal{T}$, producing the final output as:
$y=\mathcal{M}\left(\mathcal{I}_{ta},q,\mathcal{T},\mathcal{R}\right),$ (3)
where $\mathcal{I}_{ta}$ instructs the model $\mathcal{M}$ to align its reasoning with assimilated external knowledge. Through this thought accommodation process, ThinkNote enables the model to approximate cognitive disequilibrium regulation, detecting mismatches between external knowledge and internal reasoning.
3.2 Configuring ThinkNote by Mimicking Different Human Learning Behaviors
Learning behaviors are goal-directed mental operations used to enhance knowledge utilization and problem-solving. ThinkNote employs flexible instructions $\mathcal{I}_{ka}$ and $\mathcal{I}_{ta}$ to simulate diverse human learning behaviors at each stage of the cognitive process, supporting both knowledge assimilation and thought accommodation. The instruction is instantiated in one of four constructivist modes: Anchoring, Association, Reasoning, and Reflection. The details of these instruction templates are shown in Appendix A.5.
Anchoring. When humans encounter a new concept like “photosynthesis” in the query, they would start by identifying more basic concepts such as “chlorophyll” or “carbon dioxide” in external knowledge to create reference points in memory. Similarly, the instruction of anchoring behavior guides models to simulate this behavior and help them form an initial understanding by extracting related information from unfamiliar external knowledge. This behavior grounds its comprehension of queries and facilitates the expansion of topic-relevant knowledge Yu et al. (2023a).
Association. The association behavior maps new facts to known entities. Rather than introducing novel concepts, it reinforces and elaborates on familiar knowledge. For instance, when encountering information about electric vehicles, the model relates it to existing knowledge in their parametric memory of “batteries” and “charging stations”, thereby deepening its understanding by associating these new contents to previously learned concepts Steffe and Gale (1995).
Reasoning. Inspired by Edge et al. (2024), the reasoning behavior guides the model to extract and organize logical relationships found in the external knowledge. This enables LLMs to build coherent and structured explanations of external knowledge. For example, when synthesizing documents on climate change policy, the model infers causal links, such as “carbon tax leads to emission reduction, which in turn impacts long-term economic growth”—to support principled reasoning.
Reflection. This behavior encourages the model to evaluate its outputs by comparing internal knowledge with external knowledge. By treating the external documents as authoritative references, the model can detect inconsistencies and revise its responses. For instance, if the model initially generates an outdated population statistic, reflection allows it to identify and correct the discrepancy based on up-to-date knowledge Xu et al. (2024b).
4 Experimental Methodology
In this section, we describe the datasets, baselines, and implementation details of our experiments.
Dataset. Our experiment uses six datasets to evaluate the performance of ThinkNote, including NQ Kwiatkowski et al. (2019), PopQA Mallen et al. (2023), TriviaQA Joshi et al. (2017), 2WikiMultiHopQA Ho et al. (2020) and ASQA Stelmakh et al. (2022). We use Contriever Izacard et al. (2021) to obtain the external knowledge for all datasets to ensure fairness, following previous work Asai et al. (2023); Wei et al. (2024). Due to the cost of inference, we randomly sample a subset of 500 questions from each dataset in our experiments in line with prior work Trivedi et al. (2023); Yoran et al. (2023).
Method LLM PopQA NQ TriviaQA 2WikiMHQA ASQA
(acc) (acc) (acc) (acc) (rec) (str-em) (hit) (rec)
Vanilla LLMs
[…]
LLM w/ RAG
[…]
Table 1: Overall Performance of ThinkNote and Baselines on Different Knowledge-Intensive Benchmarks. For ThinkNote, we use the instruction of association learning behavior to guide both knowledge assimilation and thought accommodation and evaluate the overall performance.
Evaluation Metrics. Following Asai et al. (2023) and Wei et al. (2024), we use correctness (str-em), hit rates (hit), and recall (rec) to evaluate the question-answering performance on the ASQA task and use accuracy (acc) for other tasks. All of the evaluation metrics are implemented referring to the ALCE toolkit Gao et al. (2023a).
Baselines. The baselines consist of vanilla LLMs, LLMs with CoT, and some RAG models. We use Meta-Llama-3-Instruct and ChatGPT-4o-Mini as backbone LLMs to implement baseline models. Detailed descriptions, configurations, and implementation specifics of each baseline model are provided in Appendix A.4.
Vanilla LLMs. We employ two methods to prompt LLMs to answer questions based solely on their parametric memory. First, the Direct QA approach involves directly instructing the LLMs to answer questions. Second, we implement the Chain-of-Thought Wei et al. (2022b) prompting, which generates a step-by-step reasoning process to arrive at the answer.
RAG Models. We implement three widely used RAG models as baselines, including vanilla RAG, Chain-of-Note Yu et al. (2023a), and Self-RAG Asai et al. (2023). The vanilla RAG model directly feeds retrieved passages as the context and asks the LLM to generate the answer Ram et al. (2023a). Chain-of-Note extends the CoT method to summarize the query-related knowledge from external knowledge for producing answers. Self-RAG Asai et al. (2023) trains LLMs to employ a self-reflection mechanism to filter out irrelevant evidence effectively.
Implementation Details. We use different language models, including Meta-Llama-3-Instruct-8B, Meta-Llama-3-Instruct-70B, and ChatGPT-4o-Mini as the foundation model to evaluate the performance of the ThinkNote. Specifically, we deploy Meta-Llama-3-Instruct models using the vLLM Kwon et al. (2023), and use the OpenAI SDK to call the ChatGPT-4o-Mini API for experiments. The temperature is set to 0.2 for all models.
5 Evaluation Result
In this section, we first present the overall performance of ThinkNote. We then analyze its effectiveness in leveraging knowledge. Finally, we demonstrate how ThinkNote enhances the self-consistency of LLMs in answering questions.
5.1 Overall Performance
The overall performance of ThinkNote on knowledge-intensive tasks is shown in Table 1.
Compared to vanilla RAG models, ThinkNote delivers over a 10% improvement in performance, demonstrating its effectiveness in enabling LLMs to more accurately integrate and utilize external knowledge for question answering. While self-refinement methods focus on denoising external information to enhance generation, ThinkNote surpasses these approaches with over a 4% improvement, highlighting the importance of fostering a deeper understanding of knowledge within LLMs rather than merely extracting surface-level content related to queries. Furthermore, ThinkNote consistently outperforms all baseline models, regardless of the model scale or whether the backbone is black or white-box, indicating its strong generalization capabilities.
Compared with vanilla RAG models, the Self-Refined methods exhibit performance degradation on TriviaQA and 2WikiMHQA datasets, showing that filtering query-related contents may risk omitting critical information from external knowledge. Unlike Self-Refined RAG, ThinkNote employs a two-stage cognitive process that actively acquires essential knowledge from external information, rather than relying solely on LLMs to evaluate the relevance between queries and evidence. By leveraging constructivist cognitive modeling, ThinkNote offers a promising way to emulate human learning behaviors and demonstrates its effectiveness by consistently achieving improvements across all test scenarios.
Method Llama-3-Ins-8B Llama-3-Ins-70B
PopQA 2WikiMHQA ASQA PopQA 2WikiMHQA ASQA
(acc) (acc) (str-em) (acc) (acc) (str-em)
Cognitive Modeling Ablation
ThinkNote 65.8 53.8 43.4 69.8 63.2 48.8
w/o Knowledge Assimilation 58.4 48.2 38.6 61.8 59.6 43.2
w/o Thought Accommodation 63.0 46.8 40.2 66.0 57.4 45.7
Learning Behaviors Ablation
w. Anchoring Instruction 64.8 54.6 39.9 69.0 63.6 45.8
w. Association Instruction 65.8 53.8 43.4 69.8 63.2 48.8
w. Reasoning Instruction 64.0 52.8 40.3 68.4 59.6 48.3
w. Reflection Instruction 63.2 52.8 42.2 65.2 60.6 50.2
[…]
5.2 Ablation Study
As shown in Table 2, we conduct ablation studies to evaluate the contribution of each cognitive stage in ThinkNote. We further analyze the effectiveness of ThinkNote under different configurations of knowledge learning behavior instructions.
We first evaluate the performance of ThinkNote variants on three QA tasks: PopQA, 2WikiMHQA, and ASQA. For PopQA, the knowledge assimilation stage in ThinkNote contributes more significantly to performance improvements than the thought accommodation stage, highlighting that answering questions about less frequent entities heavily depends on acquiring information from external knowledge. In contrast, for 2WikiMHQA and ASQA, both the knowledge assimilation and thought accommodation modules are equally critical to the effectiveness of ThinkNote, emphasizing the advantages of chain-of-thought reasoning in tackling more complex QA tasks. With increasing model scale, the effectiveness of knowledge assimilation declines, while the advantages of thought accommodation remain steady. This observation implies that although larger LLMs encode more knowledge within their parameters, deliberate reasoning remains essential for effectively tackling these multi-hop QA and long-form QA tasks.
We next examine the impact of knowledge learning behaviors in ThinkNote by guiding the knowledge assimilation and thought accommodation stages using different sets of instructions, $\mathcal{I}_{ka}$ and $\mathcal{I}_{ta}$, respectively. Among all instruction types, the association learning behavior consistently achieves superior performance across diverse QA tasks. This behavior encourages LLMs to reinforce and extend their internal knowledge representations, rather than relying primarily on the assimilation of unfamiliar information from retrieved documents, as is characteristic of anchoring learning behavior. These findings emphasize the importance of integrating external knowledge through associations with the model’s parametric memory. For subsequent experiments, we adopt the association learning behavior instruction to configure ThinkNote.
(a) Information Flow Analysis for ThinkNote. We implement both Vanilla RAG and ThinkNote with Meta-Llama-3-Instruct-8B. Darker-colored points indicate higher saliency scores in information flow analysis.
5.3 Information Flow Analysis for ThinkNote in Knowledge Utilization
This section focuses on analyzing how knowledge assimilation and thought accommodation influence the reasoning process. Specifically, we randomly sample a case from the NQ dataset and follow Fan et al. (2025) to utilize the saliency score to examine interactions between tokens, which is also referred to as the information flow. Further details on the saliency score are provided in Appendix A.2.
We show the information flow for both Vanilla RAG and ThinkNote in Figure LABEL:fig:ifrag and Figure LABEL:fig:ifarag, respectively. The Vanilla RAG method demonstrates a dense and intricate pattern of information flow, where a broad range of tokens indiscriminately influence other tokens throughout the sequence. This indicates that Vanilla RAG lacks the mechanism to integrate knowledge or suppress irrelevant information, leading to an over-reliance on incidental associations. In contrast, ThinkNote guides the information flow to some specific tokens (highlighted with the green box) by modeling cognitive processes. These tokens contribute to the accumulation of information flow as part of a chain-of-thought. This indicates that ThinkNote can effectively integrate and utilize information to help LLMs generate responses, reflecting a more goal-directed reasoning process, where irrelevant or distracting information is attenuated.
We further analyze the accumulation of information flow in both Vanilla RAG and ThinkNote in Figure LABEL:fig:ss_rag and Figure LABEL:fig:ss_arag, highlighting the top 10 tokens with the highest accumulated saliency scores. In this example, the query is: “Who was the first baseball player to be featured on a postage stamp?” The correct answer is “Jackie Robinson”. For Vanilla RAG, the tokens with the highest accumulated saliency are generic terms like “stamp” or “postage”. While these words are related to the question, they are helpless in answering the question. In contrast, ThinkNote assigns higher saliency to key content words such as “jackie”, “robinson”, and “baseball”, indicating a more effective focus on the discriminative elements for problem-solving. This highlights its effectiveness in gathering useful information to enhance the knowledge utilization capability of LLMs.
(b) Performance of RAG and Retrieval Models with Top-$K$ Document Truncation. We implement both Vanilla RAG and ThinkNote with Meta-Llama-3-Instruct-8B.
Method Has-Answer Miss-Answer Internal Knowledge
PopQA NQ TriviaQA PopQA NQ TriviaQA PopQA NQ TriviaQA
Llama-3-Ins-8B
Direct QA 34.0 47.5 80.1 2.6 21.4 33.8 100.0 100.0 100.0
Vanilla RAG 85.5 76.2 91.7 3.2 6.3 18.0 89.3 79.1 88.7
Chain-of-Note 93.9 85.6 95.0 3.8 8.2 15.8 95.9 83.2 89.0
ThinkNote 90.1 80.4 97.0 8.3 25.8 36.0 99.2 92.9 97.3
Llama-3-Ins-70B
Direct QA 45.6 63.9 91.4 9.0 33.3 51.8 100.0 100.0 100.0
Vanilla RAG 89.0 83.6 95.6 7.1 12.6 22.3 84.8 82.7 87.8
Chain-of-Note 96.2 90.0 97.5 6.4 14.5 28.8 93.6 85.2 90.0
ThinkNote 95.9 88.3 97.5 12.2 33.3 54.0 93.6 94.1 98.0
Table 3: Experimental Results on Evaluating the Knowledge Usage Ability of Different RAG Models.
5.4 Effectiveness of ThinkNote in Knowledge Integration and Utilization
In this subsection, we evaluate the capability of ThinkNote to accurately utilize external knowledge to generate responses.
We evaluate the sensitivity of both the vanilla LLM and ThinkNote to noisy external knowledge by feeding them the top-$K$ retrieved documents. Figures LABEL:fig:popqars and LABEL:fig:nqrs show the top-$K$ retrieval accuracy, which reflects the concentration of query-relevant information in documents. As $K$ increases from 1 to 5, retrieval accuracy consistently improves, indicating a higher likelihood of relevant content being included. However, this improvement saturates after $K=3$, showing that additional documents tend to introduce more noise than useful information. Figures LABEL:fig:popqa and LABEL:fig:nq compare the performance of ThinkNote and baseline methods. While the performance of Vanilla RAG remains stable or even degrades when $K\geq 3$, indicating its limited robustness to noisy inputs, ThinkNote continues to benefit from the additional documents. This trend highlights ThinkNote’s superior ability to identify and utilize query-relevant information even in the presence of increased noise.
In Table 3, we then conduct three testing scenarios to evaluate the effectiveness of ThinkNote: Has-Answer, Miss-Answer, and Internal Knowledge. The Has-Answer scenario comprises queries where the external information contains the correct answers, assessing the model’s capability to effectively utilize external knowledge. Conversely, the Miss-Answer scenario involves queries where the external information fails to provide the correct answers. The Internal Knowledge scenario evaluates the ability of LLMs to resolve conflicts between internal and external knowledge.
As shown in the evaluation results, Chain-of-Note and ThinkNote methods significantly outperform Vanilla RAG, highlighting their ability to extract critical information for answering questions. In the Miss-Answer scenario, ThinkNote presents strong effectiveness by doubling the accuracy of Chain-of-Note while maintaining the performance of Direct QA. This highlights the robustness of ThinkNote in mitigating the negative impact of noise in external knowledge. Moreover, in the Internal Knowledge scenario, ThinkNote achieves the highest accuracy, illustrating its ability to alleviate conflicts between the external knowledge and the parametric memory of LLMs.
5.5 Effectiveness of ThinkNote in Improving Self-Consistency of LLMs
In this subsection, we explore the self-consistency of ThinkNote and baseline models in answering questions. Specifically, we conduct 100 sampling iterations to calculate the ratio of correction. The accuracy reflects the model’s tendency to consistently generate correct or incorrect answers. We compare three methods: Direct QA, Vanilla RAG, and ThinkNote.
We first show the accuracy of ThinkNote and baseline models in Figure LABEL:fig:consitency_acc. ThinkNote demonstrates its effectiveness by achieving a median accuracy of around 47%, outperforming the Vanilla RAG (10%) and the Direct QA (31%). Notably, the much lower median accuracy of the Vanilla RAG compared to the Direct QA suggests that retrieved documents may negatively influence LLMs’ ability to answer certain questions Xie et al. (2024). Figure LABEL:fig:consitency_density further illustrates the density distribution of answering accuracy. Compared with the Vanilla RAG, our ThinkNote improves self-consistency by reducing the density of examples with accuracy between 30% and 80%. Different from Vanilla RAG, ThinkNote increases the number of queries answered correctly, highlighting its capacity to help LLMs effectively leverage external knowledge.
(a) Answering Consistency of Different Methods. We show the performance on the 2WikiMultiHopQA dataset. All models are implemented with Meta-Llama-3-Instruct-8B.
6 Conclusion
We present ThinkNote, a cognition-inspired framework that empowers LLMs to actively construct knowledge through assimilation and accommodation, rather than passively consuming external information. By mimicking key cognitive processes from constructivist theory, ThinkNote enables more effective integration and utilization of external knowledge and improves self-consistency in correctly answering questions.
Limitation
ThinkNote demonstrates its effectiveness in integrating and utilizing external knowledge to enhance the response accuracy of white-box and black-box LLMs. However, the cognitive process modeling requires inference with the LLMs three times to generate a final answer, which brings additional time latency and increases API call costs. Furthermore, the inputs provided to the thought accommodation tend to be lengthy, as they include outputs from knowledge assimilation and chain-of-thought.
Appendix A Appendix
A.1 License
The 2WikiMultiHopQA and ASQA datasets are released under the CC BY 3.0 license, allowing users to freely use and modify the data with proper attribution. The Natural Questions (NQ) dataset is licensed under CC BY-SA 4.0, enabling users to share and modify the data, provided they attribute the source and distribute any modifications under the same license. The PopQA and TriviaQA datasets are distributed under the MIT License, which allows for free use, modification, and distribution of the data, as long as the original copyright notice is retained. All these licenses permit the datasets to be used for academic purposes.
A.2 Further Details on the Computation of Saliency Scores
We follow Fan et al. (2025) and use the saliency score to reflect the strength and direction of information flow, which refers to how much and in what way one token influences another during the reasoning process. Specifically, for each layer $l$, the saliency between tokens is defined as:
$S=\left|\sum_{h}\left(A_{h,l}\odot\frac{\partial L(x)}{\partial A_{h,l}}\right)\right|,$ (4)
where $A_{h,l}$ denotes the attention matrix of the $h$-th head in layer $l$, $L(x)$ is the loss function (cross-entropy), and $\odot$ represents the Hadamard product. This formulation combines attention weights with their gradient-based sensitivities, allowing us to capture both what the model focuses on and how crucial that focus is to the final prediction. We obtain $A_{h,l}$ from layer 21 in our experiments, using Meta-Llama-3-8B-Instruct as the foundation model.
A.3 Case Study
We choose one case in Figure 7 to show the effectiveness of ThinkNote by comparing with Vanilla RAG and Chain-of-Note models. For the question “Where was the director of film Hatchet II born?”, both Vanilla RAG and Chain-of-Note models fail to generate correct answers. Specifically, the Vanilla RAG model is misled by the retrieved documents, confusing the directors of “Hatchet II” and “Hatchet III” while answering the given question. In contrast, the Chain-of-Note model analyzes the documents in the reasoning process and generates notes that correctly identify the director as “Adam Green”. However, despite capturing this key information, the Chain-of-Note model heavily relies on extracting answers from its notes without developing a deeper understanding of external knowledge or associating it with the parametric memory of LLMs. Unlike both Vanilla RAG and Chain-of-Note methods, ThinkNote not only emphasizes the film but also identifies its director. Moreover, through using the assimilated rationale generated by the knowledge assimilation, ThinkNote accurately identifies Adam Green’s birthplace. This demonstrates ThinkNote’s strong ability to harmonize external knowledge with parametric memory, significantly improving the quality of the responses.
We then analyze the case in Figure 8. This case is from the NQ dataset and asks about the person pictured on “5000 dollar bill”. This is not a typical RAG scenario, but it happens often: the retrieved evidence only provides the background knowledge needed to answer the question while not specifying a specific answer. To generate the correct response, models need to accurately analyze the retrieval evidence and activate the parameter knowledge to conduct the reasoning process. When answering this question, Vanilla RAG is limited by its narrow interpretation of the retrieved evidence, leading it to conclude that only \$1, \$5, \$10, \$20, and \$100 dollar bills existed. The Chain-of-Note model generates brief summaries of each passage, but it struggles to provide deeper insights or draw more comprehensive analyses. In contrast, our ThinkNote model first deepens current experience based on background knowledge (Step1), then attempts to activate LLMs to use the parametric memory to continuously answer the given question (Step2), and finally accurately generates the answer based on the reasoning process (Step3).
As shown in Figure 9, another representative case highlights the limitations of LLMs in incorporating and fully understanding external knowledge. This case is from the 2WikiMultiHopQA dataset and asks for the birthplace of “Şehzade Yusuf Izzeddin’s father”. Different from the first Case, the retrieved evidence in this instance contains the crucial information required to answer the question. However, the evidence is entangled with complex logic and noise, making it challenging for the model to extract relevant information and refine the reasoning process. The Vanilla RAG model struggles to identify useful information from the retrieval evidence. Similarly, the Chain-of-Note model exhibits confusion during the summarization process: while it successfully identifies key information, it incorrectly dismisses it as irrelevant, suggesting difficulty in integrating external knowledge and resulting in a suboptimal response. In contrast, ThinkNote generates an accurate response by leveraging knowledge assimilation to accurately analyze the retrieved evidence and then refining the reasoning process through thought accommodation.
A.4 More Details of Baseline Models
In this subsection, we describe the implementation details of the baseline models in detail.
[…]
A.5 Instructions for Mimicking Different Learning Behaviors
We present all the instructions used to mimic different learning behaviors in this subsection. The anchoring behavior instructions used by ThinkNote are illustrated in Figure 11, guiding the model to ground its understanding of unfamiliar concepts by identifying essential information from external knowledge. For association behavior, ThinkNote adopts the configuration shown in Figure 12, enabling it to connect new inputs with relevant prior knowledge in its parametric memory. The reasoning behavior is supported by the instructions in Figure 13, which help the model extract and organize logical relationships for structured understanding. Finally, the reflection behavior is guided by the instructions in Figure 14, allowing the model to evaluate and revise its outputs by comparing them against reliable external sources.
Figure 10: Prompts for Baseline Models.
Figure 11: The Anchoring Instruction of ThinkNote.
Figure 12: The Association Instruction of ThinkNote.
Figure 13: The Reasoning Instruction of ThinkNote.
Figure 14: The Reflection Instruction of ThinkNote

### v4

@@
 Extensive experimental results demonstrate that ThinkNote achieves a 10% improvement over strong baseline methods on various question-answering benchmarks.
-Further analysis indicates that ThinkNote effectively integrates and utilizes external knowledge to help LLMs generate accurate responses and improves their self-consistency.
+Further analysis indicates that ThinkNote effectively integrates and utilizes external knowledge to help LLMs generate accurate responses and improve their self-consistency.
 All data and codes are available at https://github.com/OpenMatch/ThinkNote.
 ThinkNote: Enhancing Knowledge Integration and Utilization of Large Language Models via Constructivist Cognition Modeling
-Zhipeng Xu1, Zhenghao Liu1thanks: indicates corresponding author., Yukun Yan2, Shuo Wang2, Shi Yu2 Zheni Zeng2, Chaojun Xiao2, Zhiyuan Liu2, Ge Yu1 and Chenyan Xiong3 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3Language Technologies Institute, Carnegie Mellon University, United States
+Zhipeng Xu1, Zhenghao Liu1thanks: indicates corresponding author., Yukun Yan2, Shuo Wang2, Shi Yu2, Zheni Zeng2, Chaojun Xiao2, Zhiyuan Liu2, Ge Yu1 and Chenyan Xiong3 1School of Computer Science and Engineering, Northeastern University, Shenyang, China 2Department of Computer Science and Technology, Tsinghua University, Beijing, China 3Language Technologies Institute, Carnegie Mellon University, Pittsburgh, United States
 1 Introduction
 Large Language Models (LLMs), e.g.
-GPT-4 OpenAI (2023) and LLaMA Touvron et al.
+ChatGPT-4 Achiam et al.
+(2023) and LLaMA Touvron et al.
 (2023), have shown strong emergent abilities and achieved convincing performance in various NLP tasks Wei et al.
@@
 (2025).
-This limitation leads to incorrect responses and even causes degenerated responses on certain tasks Foulds et al.
+This limitation leads to incorrect responses and can even cause degraded responses on certain tasks Foulds et al.
 (2024); Shuster et al.
@@
 This two-step cognitive process is prompted to mimic different learning behaviors of humans, where we configure ThinkNote using four instructions to process external knowledge by associating familiar knowledge, anchoring unfamiliar information, extracting logical reasoning, or identifying counterfactuals.
-Our experiments in various knowledge-intensive question-answering tasks demonstrate the effectiveness of ThinkNote.
-Compared to utilizing the external knowledge directly, ThinkNote achieves a 10% improvement regardless of the scale of the LLMs used as backbone models.
-Our further analysis shows that ThinkNote causes information accumulation Fan et al.
-(2025) on specific tokens, suggesting that ThinkNote can effectively utilize critical elements of external information in generating responses.
+Our experiments on a range of knowledge-intensive question-answering tasks demonstrate the effectiveness of ThinkNote.
+Compared with directly incorporating external knowledge, ThinkNote consistently achieves around a 10% performance improvement across backbone LLMs of different scales.
+Moreover, ThinkNote exhibits stronger robustness when facing noisy, incomplete, or partially misleading external knowledge, suggesting that actively constructing and reorganizing knowledge is more reliable than passive conditioning on retrieved evidence.
+Further analysis reveals that ThinkNote leads to more concentrated information accumulation Fan et al.
+(2025) on a small set of salient tokens, indicating its ability to identify and leverage critical elements of external knowledge during response generation.
 2 Related Work
@@
 (2023b).
-These methods typically filter or reformat noisy inputs through knowledge summarization Xu et al.
+These methods typically aim to filter or reformat noisy inputs through knowledge summarization Xu et al.
 (2023), reflection Asai et al.
@@
 (2023).
+This design choice limits their ability to support adaptive knowledge integration, especially in scenarios that require the model to actively construct and revise its understanding.
 More recent efforts explore internal modifications to how LLMs represent, revise, or adapt knowledge.
@@
 (2024).
-There is still limited consensus on what constitutes effective knowledge integration at the cognitive level, and most current methods operate under engineering-driven constraints rather than principles drawn from learning theory or epistemology Banino et al.
-(2023); Floridi and Nobre (2024); Van Rooij et al.
-(2024).
+There is still limited consensus on what constitutes effective knowledge integration at the cognitive level, and most current methods operate under engineering-driven constraints rather than principles drawn from learning theory or epistemology Floridi and Nobre (2024).
 3 Methodology
@@
 3.2).
-Figure 1: The Overview of the ThinkNote Workflow.
+Figure 1: The Overview of the ThinkNote Workflow. The framework follows a two-stage process of knowledge assimilation and thought accommodation to integrate external knowledge into model reasoning.
 3.1 Enhancing Knowledge Integration and Utilization in LLMs via ThinkNote
@@
 The details of these instruction templates are shown in Appendix A.5.
+Association.
+The association behavior maps new facts to known entities.
+Rather than introducing novel concepts, it reinforces and elaborates on familiar knowledge.
+For instance, when encountering information about electric vehicles, the model relates it to existing knowledge in their parametric memory of “batteries” and “charging stations”, thereby deepening its understanding by associating these new contents to previously learned concepts Steffe and Gale (1995).
 Anchoring.
@@
 (2023a).
-Association.
-The association behavior maps new facts to known entities.
-Rather than introducing novel concepts, it reinforces and elaborates on familiar knowledge.
-For instance, when encountering information about electric vehicles, the model relates it to existing knowledge in their parametric memory of “batteries” and “charging stations”, thereby deepening its understanding by associating these new contents to previously learned concepts Steffe and Gale (1995).
 Reasoning.
@@
 The vanilla RAG model directly feeds retrieved passages as the context and asks the LLM to generate the answer Ram et al.
-(2023a).
+(2023).
 Chain-of-Note extends the CoT method to summarize the query-related knowledge from external knowledge for producing answers.
@@
 While self-refinement methods focus on denoising external information to enhance generation, ThinkNote surpasses these approaches with over a 4% improvement, highlighting the importance of fostering a deeper understanding of knowledge within LLMs rather than merely extracting surface-level content related to queries.
-Furthermore, ThinkNote consistently outperforms all baseline models, regardless of the model scale or whether the backbone is black or white-box, indicating its strong generalization capabilities.
+Furthermore, ThinkNote consistently outperforms all baseline models, regardless of the model scale or whether the backbone is black or white-box, indicating strong generalization and robustness across different model configurations.
 Compared with vanilla RAG models, the Self-Refined methods exhibit performance degradation on TriviaQA and 2WikiMHQA datasets, showing that filtering query-related contents may risk omitting critical information from external knowledge.
@@
 w. Reflection Instruction 63.2 52.8 42.2 65.2 60.6 50.2
-[…]
+Table 2: Ablation Study of ThinkNote.
+We examine the contributions of knowledge assimilation and thought accommodation, and vary instruction designs to simulate different learning behaviors.
+Experiments are conducted with Meta-Llama-3-Instruct-8B and Meta-Llama-3-Instruct-70B.
 5.2 Ablation Study
@@
 5.3 Information Flow Analysis for ThinkNote in Knowledge Utilization
-This section focuses on analyzing how knowledge assimilation and thought accommodation influence the reasoning process.
-Specifically, we randomly sample a case from the NQ dataset and follow Fan et al.
-(2025) to utilize the saliency score to examine interactions between tokens, which is also referred to as the information flow.
+We further analyze how knowledge assimilation and thought accommodation influence the reasoning process in Figure 5(a).
+Specifically, we sample a case from the NQ dataset and follow Fan et al.
+(2025) in utilizing the saliency score to examine interactions between tokens, which is also referred to as information flow.
 Further details on the saliency score are provided in Appendix A.2.
@@
 Table 3: Experimental Results on Evaluating the Knowledge Usage Ability of Different RAG Models.
+Results are reported under three scenarios, evaluating the capabilities of models in external knowledge utilization, robustness to missing evidence, and knowledge conflict resolution.
 5.4 Effectiveness of ThinkNote in Knowledge Integration and Utilization
-In this subsection, we evaluate the capability of ThinkNote to accurately utilize external knowledge to generate responses.
-We evaluate the sensitivity of both the vanilla LLM and ThinkNote to noisy external knowledge by feeding them the top-$K$ retrieved documents.
+In this subsection, we evaluate the capability of ThinkNote to utilize external knowledge to generate responses accurately.
+As shown in Figure 5(b), we evaluate the sensitivity of both the vanilla RAG and ThinkNote to noisy external knowledge by feeding them the top-$K$ retrieved documents.
 Figures LABEL:fig:popqars and LABEL:fig:nqrs show the top-$K$ retrieval accuracy, which reflects the concentration of query-relevant information in documents.
@@
 5.5 Effectiveness of ThinkNote in Improving Self-Consistency of LLMs
-In this subsection, we explore the self-consistency of ThinkNote and baseline models in answering questions.
+In this subsection, we explore the self-consistency of ThinkNote and baseline models in answering questions, as shown in Figure 7(a).
 Specifically, we conduct 100 sampling iterations to calculate the ratio of correction.
-The accuracy reflects the model’s tendency to consistently generate correct or incorrect answers.
+The accuracy reflects the model’s tendency to generate correct or incorrect answers consistently.
 We compare three methods: Direct QA, Vanilla RAG, and ThinkNote.
+For Vanilla RAG and Direct QA, the answer generation module is used for sampling, while ThinkNote utilizes the thought accommodation to generate sampled responses.
+(a) Answering Consistency of Different Methods. We show the performance on the 2WikiMultiHopQA dataset. All models are implemented with Meta-Llama-3-Instruct-8B.
 We first show the accuracy of ThinkNote and baseline models in Figure LABEL:fig:consitency_acc.
 ThinkNote demonstrates its effectiveness by achieving a median accuracy of around 47%, outperforming the Vanilla RAG (10%) and the Direct QA (31%).
-Notably, the much lower median accuracy of the Vanilla RAG compared to the Direct QA suggests that retrieved documents may negatively influence LLMs’ ability to answer certain questions Xie et al.
+Notably, the substantially lower median accuracy of Vanilla RAG compared to Direct QA indicates that incorporating retrieved documents can hinder the LLM’s ability to produce correct answers, rather than improve it Xie et al.
 (2024).
@@
 Compared with the Vanilla RAG, our ThinkNote improves self-consistency by reducing the density of examples with accuracy between 30% and 80%.
-Different from Vanilla RAG, ThinkNote increases the number of queries answered correctly, highlighting its capacity to help LLMs effectively leverage external knowledge.
-(a) Answering Consistency of Different Methods. We show the performance on the 2WikiMultiHopQA dataset. All models are implemented with Meta-Llama-3-Instruct-8B.
+In contrast to vanilla RAG, ThinkNote substantially increases the number of queries answered correctly, highlighting its stronger capability to guide LLMs in effectively leveraging external knowledge to produce reliable answers.
 6 Conclusion
@@
 Furthermore, the inputs provided to the thought accommodation tend to be lengthy, as they include outputs from knowledge assimilation and chain-of-thought.
+Acknowledgments
+This work is partly supported by the National Natural Science Foundation of China (No. 62576082 and No. 62461146205) and CCF-Zhipu Large Model Innovation Fund (No. 202403).
 Appendix A Appendix
@@
 This formulation combines attention weights with their gradient-based sensitivities, allowing us to capture both what the model focuses on and how crucial that focus is to the final prediction.
-We obtain $A_{h,l}$ from layer 21 in our experiments, using Meta-Llama-3-8B-Instruct as the foundation model.
-A.3 Case Study
+We extract $A_{h,l}$ from the 21st layer, using Meta-Llama-3-8B-Instruct as the foundation model.
+A.3 More Details of Baseline Models
+In this subsection, we describe the implementation details of the baseline models in detail.
+[…]
+A.4 Case Study
 We choose one case in Figure 7 to show the effectiveness of ThinkNote by comparing with Vanilla RAG and Chain-of-Note models.
@@
 In contrast, ThinkNote generates an accurate response by leveraging knowledge assimilation to accurately analyze the retrieved evidence and then refining the reasoning process through thought accommodation.
-A.4 More Details of Baseline Models
-In this subsection, we describe the implementation details of the baseline models in detail.
-[…]
 A.5 Instructions for Mimicking Different Learning Behaviors